OpenStudy (anonymous):

X square minus 4x plus 4 divided by 2x square minus eight

OpenStudy (anonymous):

X^2 -4x +4/2x^2-8

OpenStudy (anonymous):

oh yeah plz telll me the solution i knw bt a littl bit confused

OpenStudy (anonymous):

hmmm x^2-4x+4/2x^2-8=0 3x^2 -4x =-12 9x-4x=-12 5x=-12 something like that?